OpenStudy (anonymous):

Suppose the area of a square is 64. What is its side length?

OpenStudy (anonymous):

Area of square = side^2

hero (hero):

\[A = s^2\]\[64 = s^2\]8 = s

OpenStudy (anonymous):

a^2=64; a=sqrt(64) a=8